OpenStudy (anonymous):

\[\lim_{x \to 0}\frac{x+1}{x^{2}(x+7)}\]

terenzreignz (terenzreignz):

Top goes to 1 bottom goes to 0....

OpenStudy (anonymous):

Thank you. I sorta noticed that myself. The answer is \[\infty \] but I want to know the middle part

OpenStudy (anonymous):

Yo can should it grows with \(\displaystyle\frac{1}{x^2}\) so their ratio when x->0 is 1.

OpenStudy (anonymous):

Ok figured it out. If you evaluate the expression to the left and right of 0, that is minute positive and negative numbers, you can see that it goes to \[ \infty \]

OpenStudy (anonymous):

Great job!